3.6.18 Draw a second spring and mass hanging from the first. If the masses are mq,m2 and the spring constants are C1,C2, theWould appreciate any help. Will upvote.
Answer #1

Ho k+P = I piatto. Pe? + < CUP to C2(02-01). 2m2 m2 consider, ƏH/39,- tipit lo) - Pl/m, = alle and = str. Apz = P2 at ap2 m2Four Hamilton equation a au2 at dp = Pz ma = CaCU2-0,7 - GIUL dt . dP - Cocom-02) consider consider, dus efet pas me dpn C2 (